OpenStudy (bananas1234):

8x^2y^2 - 4x^2y - 12xy -------------------------- 4xy

OpenStudy (anonymous):

do you just want to simplify it?

OpenStudy (anonymous):

4xy[(2xy)-x-3)]/4xy cancel the 4xy, and you get 2xy-x-3

OpenStudy (michele_laino):

we can factor out 4xy at numerator, so we can write this: \[\Large \frac{{8{x^2}{y^2} - 4{x^2}y - 12xy}}{{4xy}} = \frac{{4xy\left( {2xy - x - 3} \right)}}{{4xy}} = ...?\]